OpenStudy (anonymous):

the segment DE has a midpoint M(3,-5/2)and the endpoint D(4, -2). Find the coordinates of E

OpenStudy (anonymous):

So the segment is a straight line and you can find the slope: m = ((-5/2) - (-2)) / (3 - 4) m = 0.5 The midpoint is half way between the points D and E. Therefore the distance between the midpoint and D is the same distance as between the midpoint and E: d = sqrt( (4-3)^2 + (-2-(-5/2))^2 ) d =~1.12 Does that help? You can make an equation for the line and see where the point on the line is at a distance of 1.12 from the midpoint.

OpenStudy (anonymous):

the answer choices are (-7,10), (2,-3), (10,-7), (-3,2)

OpenStudy (anonymous):

Have you tried it yourself? How far can you get?